# From Noise to Laws: Regularized Time-Series Forecasting via Denoised Dynamic Graphs

### v1

@@ Abstract
Long-horizon multivariate time-series forecasting is challenging because realistic predictions must (i) denoise heterogeneous signals, (ii) track time-varying cross-series dependencies, and (iii) remain stable and physically plausible over long rollout horizons. We present PRISM, which couples a score-based diffusion preconditioner with a dynamic, correlation-thresholded graph encoder and a forecast head regularized by generic physics penalties. We prove contraction of the induced horizon dynamics under mild conditions and derive Lipschitz bounds for graph blocks, explaining the model’s robustness. On six standard benchmarks (Electricity, Traffic, Weather, ILI, Exchange Rate, ETT), PRISM achieves consistent SOTA with strong MSE and MAE gains. Frequency-domain analysis shows preserved fundamentals and attenuated high-frequency noise, while ablations attribute improvements to (i) denoise-aware topology, (ii) adaptivity of the graph, (iii) reaction–diffusion stabilization, and (iv) tail control via kinematic constraints. Together, these results indicate that denoising, dynamic relational reasoning, and physics-aware regularization are complementary and necessary for reliable long-horizon forecasting.
@@ 1 Introduction
Long-horizon multivariate time-series forecasting (LTSF) remains challenging because models must simultaneously (i) denoise and robustly encode local/meso-scale patterns under domain-specific noise, (ii) capture evolving cross-series interactions that are often sparse and time-varying, and (iii) respect physical regularities so that predictions remain plausible and interpretable beyond the training distribution. Deep learning models capture time-series patterns with well-designed architectures spanning a wide range of foundational backbones, including CNNs (Wang et al., 2023; Wu et al., 2023a; Hewage et al., 2020), RNNs (Lai et al., 2018; Qin et al., 2017; Salinas et al., 2020), Transformers (Vaswani et al., 2017) and MLPs (Zeng et al., 2023a; Zhang et al., 2022; Oreshkin et al., 2019; Challu et al., 2023). Transformer variants have since pushed sequence modeling forward, but their raw self-attention often underperforms or becomes brittle under long horizons and distribution shifts in LTSF benchmarks (Vaswani et al., 2017; Zhou et al., 2021a; Wu et al., 2021; Zhou et al., 2022; Nie et al., 2023; Wu et al., 2023b; Liu et al., 2024; Zeng et al., 2023b). In parallel, graph neural networks (GNNs) excel at encoding relational inductive biases over sensor networks and multivariate channels, yet most approaches assume static or weakly-adaptive graphs and struggle to integrate uncertainty-aware denoising with interpretable constraints (Li et al., 2018; Yu et al., 2018; Wu et al., 2019; 2020). Diffusion generative models offer strong denoising priors, particularly when signals are corrupted or partially observed, but they are rarely tightly coupled with forecasting architectures and physical regularization in a single, end-to-end pipeline (Song et al., 2021; Ho et al., 2020; Tashiro et al., 2021). These gaps motivate our design.
We propose PRISM, a denoised and physics-regularized inter-series structure model that (a) preconditions input series through a score-based diffusion denoiser to recover fine-scale structure before feature extraction; (b) constructs a dynamic, functionally linked graph whose edges are induced by data-driven inter-series dependence and evolve with time, enabling bidirectional message passing among series and across temporal features; and (c) injects domain-agnostic and physics-informed constraints during training, yielding interpretable forecasts that satisfy generic conservation/smoothness priors without hard-crafting task-specific equations (Raissi et al., 2019; Karniadakis et al., 2021; Shuman et al., 2013; Dong et al., 2019). The result is a single, coherent model that unifies uncertainty-aware denoising, dynamic relational reasoning, and physically grounded regularization. Our contributions are summarized as follows:
  • Diffusion as a preconditioner for LTSF. We apply score-based diffusion to the history before forecasting to suppress noise and accentuate coherent modes—rather than generating sequences post hoc—yielding robustness to covariate shift and low-SNR regimes (Song et al., 2021; Ho et al., 2020; Tashiro et al., 2021).
  • Dynamic correlation-thresholded, function-linked graphs. Sliding-window correlations and functional couplings define a time-varying graph where edges appear only above data-driven thresholds, producing sparse, interpretable topologies with bidirectional spatio–temporal message passing (Li et al., 2018; Yu et al., 2018; Wu et al., 2019; 2020; Shuman et al., 2013; Dong et al., 2019).
  • Physics-informed regularization for interpretability and stability. Generic physics-motivated soft constraints (smoothness, bounded variation, energy/dissipation surrogates) in the forecast head promote physically plausible rollouts and clearer attributions without requiring domain-specific PDEs (Raissi et al., 2019; Karniadakis et al., 2021).
Beyond empirical accuracy, we provide theoretical results (see Methodology) establishing identifiability of the denoising-plus-forecasting objective under mild conditions and the stability of message passing on the dynamically thresholded graph, clarifying why the three ingredients work better together than in isolation. Together, these contributions directly address the shortcomings of prior Transformer-only, GNN-only, or diffusion-only pipelines on standard LTSF benchmarks.
@@ 2 Related Works
Early progress in sequence modeling was driven by the Transformer (Vaswani et al., 2017), inspiring LTSF variants that capture long-range dependencies more efficiently, Informer with ProbSparse attention (Zhou et al., 2021a), Autoformer with trend/seasonal decomposition and auto-correlation (Wu et al., 2021), and FEDformer via frequency-domain modeling (Zhou et al., 2022). Newer designs, PatchTST (patching, channel independence) (Nie et al., 2023), TimesNet (2D temporal variations) (Wu et al., 2023b), and iTransformer (axis inversion to emphasize variate tokens) (Liu et al., 2024), further reduce complexity and exploit multivariate structure. Yet DLinear and the LTSF-Linear family show that, on common benchmarks, simple linear forecasters can rival or outperform many transformers, challenging whether permutation-invariant self-attention aligns with ordered temporal dynamics for long horizons (Zeng et al., 2023b). Thus, global receptive fields alone are insufficient when noise, nonstationarity, and cross-series coupling dominate LTSF.
@@
Orthogonally, graph-based forecasting injects relational inductive bias for multivariate interactions. DCRNN models diffusion on road networks, STGCN alternates graph and temporal convolutions, Graph WaveNet learns adaptive adjacency via node embeddings, and MTGNN jointly learns directed graphs and temporal convolutions (Li et al., 2018; Yu et al., 2018; Wu et al., 2019; 2020). These works show that who influences whom matters as much as temporal depth. Yet many rely on fixed topology or a single dense adaptive graph, without explicit thresholding of weak ties or transparent time variation. Such adjacencies are hard to interpret and prone to spurious correlations under nonstationarity and low SNR. We instead construct dynamic, correlation-thresholded graphs: retaining edges only when dependence (or functional coupling) exceeds a principled threshold yields sparse, interpretable and bidirectional topologies, which are consistent with correlation-network practice (e.g., MST/PMFG) for revealing hierarchical structure (Shuman et al., 2013; Dong et al., 2019).
@@
On the uncertainty and denoising side, diffusion probabilistic models and score-based SDEs have established new generative baselines with principled noise injection and reverse-time denoising (Song et al., 2021; Ho et al., 2020). In time-series, CSDI adapts score-based diffusion for conditional imputation across channels and time, demonstrating robustness to missingness and noise (Tashiro et al., 2021). Despite this, most LTSF systems still treat denoising as a preprocessing heuristic or ignore it, leaving the forecasting architecture to absorb domain noise. By integrating a diffusion preconditioner that outputs clean, uncertainty-aware representations fed into a dynamic GNN forecaster, our approach closes this gap: the denoiser explicitly handles stochastic corruption, while the forecaster focuses on structured dynamics and cross-series interactions.
@@
Finally, physics-informed neural networks (PINNs) and related physics-guided regularization inject inductive biases via soft penalties derived from differential operators or conservation laws, promoting data efficiency and interpretability (Raissi et al., 2019; Karniadakis et al., 2021). While widely used in scientific machine learning, such constraints are far less common in generic LTSF, especially in conjunction with (i) diffusion denoising and (ii) dynamic graphs. Our design adopts domain-agnostic physics surrogates (e.g., smoothness/energy/monotonicity budgets) that are meaningful across diverse LTSF datasets (electricity load, traffic occupancy, meteorology, epidemiology, exchange rates, and transformer telemetry) (Zhou et al., 2021b; Repository, 2014; PEMS-SF, 2017; Lai, 2017; for Disease Control & Prevention, 2021; Rasul et al., 2024), delivering (a) calibrated, physically plausible trajectories without brittle hard constraints and (b) interpretable attributions via constraint-specific penalties.
@@
In summary, prior Transformers emphasize long-range token mixing but are fragile under noise and cross-series nonstationarity (Zhou et al., 2021a; Wu et al., 2021; Zhou et al., 2022; Nie et al., 2023; Wu et al., 2023b; Liu et al., 2024; Zeng et al., 2023b); graph forecasters encode relations but often with static or opaque connectivity (Li et al., 2018; Yu et al., 2018; Wu et al., 2019; 2020); and diffusion or physics-guided components are seldom coupled tightly with forecasting to address denoising and plausibility together. PRISM is necessary—not a “stitching” of fashionable modules—because each component resolves a distinct, documented deficiency and the pipeline is co-designed: diffusion improves SNR for graph reasoning; dynamic, thresholded graphs expose interpretable dependencies for message passing; and physics-informed penalties regularize the forecast trajectory where pure data fitting over-extrapolates. The overall architecture of DORIC is illustrated in Figure 1 .
@@ 3.1 Problem Setup and Notation
$\begin{aligned} &\underbrace{x_{t-L+1:t}=X[t-L+1:t]\in\mathbb{R}^{L\times N}}_{\text{history}},\\[3.0pt] &\underbrace{y_{t+1:t+H}=X[t+1:t+H]\in\mathbb{R}^{H\times N}}_{\text{future}},\end{aligned}\qquad\smash{\vbox{\hbox{$t=L,\ldots,T-H$}}}.$
@@ 3.2 Series-wise Denoising via Diffusion Model
$\mathcal{L}_{\mathrm{ddpm}}=\mathbb{E}_{x_{0},\varepsilon,t}\big\|\varepsilon-\varepsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon,\,t)\big\|_{2}^{2}.$
@@ 3.3 Dynamic Graph Construction from Correlations
$C_{t}(i,j)=\mathrm{Corr}\!\left(x_{t-W+1:t,i},\,x_{t-W+1:t,j}\right).$
@@
$\bar{A}_{t}\;=\;D_{t}^{-\frac{1}{2}}(A_{t}+I)D_{t}^{-\frac{1}{2}},\qquad D_{t}=\mathrm{diag}((A_{t}+I)\mathbbm{1}),$
@@ 3.6 Configurable Decoder
$\hat{y}_{t+1:t+H}\;=\;\Psi\!\left(H^{(L_{g})}_{t}\right)\in\mathbb{R}^{H\times D}.$
@@ 3.7 Physics- and Structure-Aware Regularizers
$\mathcal{L}_{\mathrm{data}}\;=\;\frac{1}{DH}\sum_{i=1}^{D}\sum_{h=1}^{H}\Big(\hat{y}_{h,i}-y_{h,i}\Big)^{2}.$
@@
$\mathcal{L}_{\mathrm{range}}=\frac{1}{DH}\sum_{i=1}^{D}\sum_{h=1}^{H}\left(\,[m_{i}-\hat{y}_{h,i}]_{+}^{2}+[\hat{y}_{h,i}-M_{i}]_{+}^{2}\right).$
@@ 3.8 Theoretical Properties
We present two propositions that explain stability and regularity of PRISM under mild conditions encountered in practice(proof details in the Appendix  C).
@@ 4.1 Experimental setting & baselines
Model Electricity Traffic Weather ILI Exchange Rate ETT
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
LogTrans 0.272 0.370 0.705 0.395 0.696 0.602 4.480 1.444 0.968 0.812 1.534 0.899
Informer 0.311 0.397 0.764 0.416 0.634 0.548 5.764 1.677 0.847 0.752 1.410 0.810
Autoformer 0.227 0.338 0.628 0.379 0.338 0.382 3.483 1.287 0.197 0.323 0.327 0.371
FEDformer 0.214 0.327 0.610 0.376 0.309 0.360 2.203 0.963 0.183 0.297 0.305 0.349
Crossformer 0.244 0.334 0.667 0.426 0.264 0.320 1.572 0.891 0.175 0.293 0.757 0.610
TimesNet 0.193 0.304 0.620 0.336 0.251 0.294 1.365 0.806 0.158 0.281 0.291 0.333
PatchTST 0.216 0.318 0.529 0.341 0.265 0.285 0.952 0.793 0.146 0.276 0.290 0.334
TimeMixer 0.182 0.272 0.484 0.297 0.240 0.271 0.877 0.763 0.117 0.258 0.275 0.323
PRISM 0.156 0.228 0.375 0.218 0.211 0.239 0.672 0.505 0.088 0.196 0.258 0.291
Table 1: Results on six benchmarks. The results on Electricity , Traffic, Weather and ETT are averaged from 4 different prediction lengths, that is [96,192,336,720]. The results on ILI are from 24 prediction length and the results on Exchange Rate are from 96 prediction length.

### v2

# Diffusion Denoised and Physics-regularized Inter-series Model for Long-horizon Multivariate Time-series Forecasting

@@ Abstract
Problem and application. Long-horizon multivariate time-series forecasting (LTSF) underpins engineering decision-making in electricity demand, road traffic, weather sensing, influenza monitoring, and exchange-rate analytics. Accurate forecasts should simultaneously denoise heterogeneous sensor streams, track time-varying cross-series couplings, and remain stable and physically plausible over long rollouts.
AI contribution. We present PRISM (a denoised and Physics-Regularized Inter-Series Model) that causally links three components into an organic whole: a score-based diffusion preconditioner raises the effective signal-to-noise ratio of histories; on the cleaned signals, a correlation-thresholded dynamic graph encodes evolving inter-series dependence with sparsity for interpretability; and a reaction–diffusion regularizer in the forecast head provides physics-inspired (structure-aware) stabilization of multi-step rollouts. This chain implements the hypothesis denoise $\Rightarrow$ estimate reliable sparse dynamics $\Rightarrow$ stabilize, which we motivate theoretically and validate empirically.
Results and evidence. Under standardized tuning budgets and reporting of wall-clock cost, PRISM yields consistent improvements in MSE/MAE across six established benchmarks (Electricity, Traffic, Weather, ILI, Exchange Rate, ETT). Frequency-domain diagnostics indicate preserved fundamentals with attenuated spurious high-frequency bursts, and ablations attribute gains to (i) denoise-aware topology, (ii) adaptivity and sparsity of the graph, and (iii) reaction–diffusion stabilization with simple kinematic range priors. We do not introduce new physical laws; constraints are domain-agnostic and serve as regularizers. Together, these results suggest that uncertainty-aware denoising, dynamic relational reasoning, and physics-inspired stabilization are complementary and necessary for reliable long-horizon engineering forecasts.
Authors:
Hongwei Maa, Jiayu Fanga, Dai Shib, Minh-Ngoc Trana, Junbin Gaoa,∗
Affiliations:
a
The University of Sydney, City Rd, Darlington, NSW 2006, Sydney, Australia
b The University of Cambridge, Trumpington Street, Cambridge CB2 1PZ, Cambridge, UK
Keywords: Time-series forecasting; Diffusion Denoising; Dynamic Correlation Graphs; Physics Regularization; Reaction-diffusion Stabilization.
@@ 1 Introduction
Background and challenge.
Long-horizon multivariate time-series forecasting (LTSF) is central to many engineering systems, from power grids and transportation networks to meteorological and public-health monitoring. Forecasts must meet three coupled requirements: (i) denoise and robustly encode local/meso-scale patterns under domain-specific disturbances; (ii) capture evolving cross-series interactions that are often sparse and regime-dependent; and (iii) respect generic physical regularities so that multi-step rollouts remain plausible and interpretable beyond the training distribution.
Gaps in existing paradigms.
Deep learning backbones—CNNs (32; 33; 34), RNNs (35; 36; 37), Transformers (1), and MLPs (38; 39; 40; 41)—have advanced sequence modeling, yet on LTSF benchmarks the raw self-attention stack can be brittle at long horizons and under shifts (1; 2; 3; 4; 5; 6; 7; 8). In parallel, graph neural networks (GNNs) encode relational inductive biases for sensor arrays and multivariate channels, but many methods assume static or weakly adaptive graphs and rarely combine uncertainty-aware denoising with interpretable constraints (13; 14; 15; 16). Diffusion models provide powerful denoising priors, especially with partial observation or low SNR, yet they are seldom tightly integrated with forecasting architectures and physics-oriented regularization in a single end-to-end pipeline (17; 18; 19). These gaps motivate our design.
Design hypothesis: a causal chain.
We posit that three inexpensive priors should be composed in order:
  1. 1.
    Diffusion preconditioning increases effective SNR in the input window, shrinking spurious high-frequency bursts that otherwise inflate false correlations.
  2. 2.
    Correlation-thresholded dynamic graphs estimated on the cleaned signals then track regime-dependent couplings with explicit sparsity, improving interpretability and reducing error propagation across weakly related series.
  3. 3.
    Reaction–diffusion stabilization in the forecast head penalizes spatially incoherent activations (diffusion term) while allowing task-driven adjustments (reaction term), damping multi-step drift without hard-coding domain-specific PDEs.
The chain implements denoise $\Rightarrow$ estimate dynamics $\Rightarrow$ stabilize. Breaking the order re-introduces specific pathologies: without 1, the graph misconnects under noise; without 2, stale edges hinder regime shifts; without 3, rollouts amplify residual bias. Our theoretical analysis provides contraction of the induced horizon dynamics under mild conditions and Lipschitz bounds for the graph blocks, clarifying why the composition improves robustness.
The PRISM framework.
PRISM operationalizes this hypothesis by (a) applying score-based diffusion to the history before feature extraction; (b) constructing a sliding-window, correlation-thresholded, function-linked graph with bidirectional spatio–temporal message passing (13; 14; 15; 16; 22; 23); and (c) injecting domain-agnostic, physics-inspired soft constraints (smoothness, bounded variation, simple energy/dissipation surrogates) during training (20; 21; 22; 23). We stress that we do not propose new physical laws; the constraints act as structure-aware regularizers that improve stability and attribution.
Contributions.
  • A unified, causal pipeline for LTSF. We couple diffusion preconditioning, dynamic sparse graphs, and reaction–diffusion stabilization into a single end-to-end model, turning common heuristics into a checkable design with explicit failure-mode analyses.
  • Theory for robustness. We establish contraction of horizon dynamics under mild assumptions and derive Lipschitz bounds for the dynamically thresholded graph blocks, explaining when and why the composition is stable.
  • Transparent evidence under standardized budgets. On six benchmarks (Electricity, Traffic, Weather, ILI, Exchange Rate, ETT), we report mean/variance metrics, per-horizon breakdowns, sensitivity to thresholds/weights, and wall-clock costs. Frequency-domain diagnostics show preserved fundamentals with attenuated spurious high-frequency components.
Positioning.
Compared to Transformer-only stacks that may overfit noise (1; 2; 3; 4; 5; 6; 7; 8), PRISM first raises SNR; compared to GNN-only pipelines that assume fixed topology (13; 14; 15; 16), PRISM adapts edges in time; and compared to diffusion-only approaches (17; 18; 19), PRISM regularizes rollouts with physics-inspired stability priors. This unified view addresses accuracy, interpretability, and stability in engineering LTSF without relying on task-specific PDEs.
@@ 2 Related Works
Early progress in sequence modeling was driven by the Transformer (1), inspiring LTSF variants that capture long-range dependencies more efficiently, Informer with ProbSparse attention (2), Autoformer with trend/seasonal decomposition and auto-correlation (3), and FEDformer via frequency-domain modeling (4). Newer designs, PatchTST (patching, channel independence) (5), TimesNet (2D temporal variations) (6), and iTransformer (axis inversion to emphasize variate tokens) (7), further reduce complexity and exploit multivariate structure. Yet DLinear and the LTSF-Linear family show that, on common benchmarks, simple linear forecasters can rival or outperform many transformers, challenging whether permutation-invariant self-attention aligns with ordered temporal dynamics for long horizons (8). Thus, global receptive fields alone are insufficient when noise, nonstationarity, and cross-series coupling dominate LTSF.
@@
Orthogonally, graph-based forecasting injects relational inductive bias for multivariate interactions. DCRNN models diffusion on road networks, STGCN alternates graph and temporal convolutions, Graph WaveNet learns adaptive adjacency via node embeddings, and MTGNN jointly learns directed graphs and temporal convolutions (13; 14; 15; 16). These works show that who influences whom matters as much as temporal depth. Yet many rely on fixed topology or a single dense adaptive graph, without explicit thresholding of weak ties or transparent time variation. Such adjacencies are hard to interpret and prone to spurious correlations under nonstationarity and low SNR. We instead construct dynamic, correlation-thresholded graphs: retaining edges only when dependence (or functional coupling) exceeds a principled threshold yields sparse, interpretable and bidirectional topologies, which are consistent with correlation-network practice (e.g., MST/PMFG) for revealing hierarchical structure (22; 23).
@@
On the uncertainty and denoising side, diffusion probabilistic models and score-based SDEs have established new generative baselines with principled noise injection and reverse-time denoising (17; 18). In time-series, CSDI adapts score-based diffusion for conditional imputation across channels and time, demonstrating robustness to missingness and noise (19). Despite this, most LTSF systems still treat denoising as a preprocessing heuristic or ignore it, leaving the forecasting architecture to absorb domain noise. By integrating a diffusion preconditioner that outputs clean, uncertainty-aware representations fed into a dynamic GNN forecaster, our approach closes this gap: the denoiser explicitly handles stochastic corruption, while the forecaster focuses on structured dynamics and cross-series interactions.
@@
Finally, physics-informed neural networks (PINNs) and related physics-guided regularization inject inductive biases via soft penalties derived from differential operators or conservation laws, promoting data efficiency and interpretability (20; 21). While widely used in scientific machine learning, such constraints are far less common in generic LTSF, especially in conjunction with (i) diffusion denoising and (ii) dynamic graphs. Our design adopts domain-agnostic physics surrogates (e.g., smoothness/energy/monotonicity budgets) that are meaningful across diverse LTSF datasets (electricity load, traffic occupancy, meteorology, epidemiology, exchange rates, and transformer telemetry) (26; 27; 28; 29; 30; 31), delivering (a) calibrated, physically plausible trajectories without brittle hard constraints and (b) interpretable attributions via constraint-specific penalties.
@@
In summary, prior Transformers emphasize long-range token mixing but are fragile under noise and cross-series nonstationarity (2; 3; 4; 5; 6; 7; 8); graph forecasters encode relations but often with static or opaque connectivity (13; 14; 15; 16); and diffusion or physics-guided components are seldom coupled tightly with forecasting to address denoising and plausibility together. PRISM is necessary because each component resolves a distinct, documented deficiency and the pipeline is co-designed: diffusion improves SNR for graph reasoning; dynamic, thresholded graphs expose interpretable dependencies for message passing; and physics-informed penalties regularize the forecast trajectory where pure data fitting over-extrapolates. The overall architecture of DORIC is illustrated in Figure 1 .
@@ 3.1 Problem Setup and Notation
$\begin{aligned} &\underbrace{x_{t-L+1:t}=X[t-L+1:t]\in\mathbb{R}^{L\times N}}_{\text{history}},\\[3.0pt] &\underbrace{y_{t+1:t+H}=X[t+1:t+H]\in\mathbb{R}^{H\times N}}_{\text{future}},\end{aligned}\qquad\smash{\vbox{\hbox{$t=L,\ldots,T-H$}}}.$ (1)
@@ 3.2 Series-wise Denoising via Diffusion Model
$\mathcal{L}_{\mathrm{ddpm}}=\mathbb{E}_{x_{0},\varepsilon,t}\big\|\varepsilon-\varepsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon,\,t)\big\|_{2}^{2}.$ (2)
@@ 3.3 Dynamic Graph Construction from Correlations
$C_{t}(i,j)=\mathrm{Corr}\!\left(x_{t-W+1:t,i},\,x_{t-W+1:t,j}\right).$ (3)
@@
$A_{t}(i,j)=\mathbf{1}\!\left(|C_{t}(i,j)|>\tau\right)\cdot|C_{t}(i,j)|^{\gamma},\qquad A_{t}(i,i)=0,$ (4)
@@
$\bar{A}_{t}\;=\;D_{t}^{-\frac{1}{2}}(A_{t}+I)D_{t}^{-\frac{1}{2}},\qquad D_{t}=\mathrm{diag}((A_{t}+I)\mathbbm{1}),$ (5)
@@ 3.6 Configurable Decoder
$\hat{y}_{t+1:t+H}\;=\;\Psi\!\left(H^{(L_{g})}_{t}\right)\in\mathbb{R}^{H\times D}.$ (9)
@@ 3.7 Physics- and Structure-Aware Regularizers
$\mathcal{L}_{\mathrm{data}}\;=\;\frac{1}{DH}\sum_{i=1}^{D}\sum_{h=1}^{H}\Big(\hat{y}_{h,i}-y_{h,i}\Big)^{2}.$ (10)
@@
$\mathcal{L}_{\mathrm{range}}=\frac{1}{DH}\sum_{i=1}^{D}\sum_{h=1}^{H}\left(\,[m_{i}-\hat{y}_{h,i}]_{+}^{2}+[\hat{y}_{h,i}-M_{i}]_{+}^{2}\right).$ (11)
@@ 3.8 Theoretical Properties
We present two propositions that explain stability and regularity of PRISM under mild conditions encountered in practice(proof details in the Appendix  B).
@@
Why it stabilizes long-horizon rollouts.
Proposition 1 ensures $\rho(M)<1$ for $0<\kappa,\gamma$ with $\kappa+\gamma<1$, giving temporal contraction; Proposition 2 bounds the Lipschitz constant of graph blocks, keeping spatial mixing well-conditioned. This explains the empirical robustness of the full composition (denoise $\Rightarrow$ dynamics $\Rightarrow$ stabilize).
@@ 4.1 Experimental setting & baselines
Table 1: Results on six benchmarks. Entries are mean $\pm$ std over 10 random seeds.
LogTrans Informer Autoformer FEDformer Crossformer TimesNet PatchTST TimeMixer PRISM
Electricity
MSE $0.272\pm 0.009$ $0.311\pm 0.010$ $0.227\pm 0.005$ $0.214\pm 0.004$ $0.244\pm 0.005$ $0.193\pm 0.004$ $0.216\pm 0.004$ $0.182\pm 0.003$ $0.156\pm 0.003$
MAE $0.370\pm 0.006$ $0.397\pm 0.008$ $0.338\pm 0.004$ $0.327\pm 0.004$ $0.334\pm 0.004$ $0.304\pm 0.004$ $0.318\pm 0.003$ $0.272\pm 0.003$ $0.228\pm 0.003$
Traffic
MSE $0.705\pm 0.015$ $0.764\pm 0.018$ $0.628\pm 0.011$ $0.610\pm 0.010$ $0.667\pm 0.013$ $0.620\pm 0.009$ $0.529\pm 0.007$ $0.484\pm 0.006$ $0.375\pm 0.006$
MAE $0.395\pm 0.010$ $0.416\pm 0.012$ $0.379\pm 0.009$ $0.376\pm 0.008$ $0.426\pm 0.012$ $0.336\pm 0.007$ $0.341\pm 0.004$ $0.297\pm 0.003$ $0.218\pm 0.004$
Weather
MSE $0.696\pm 0.020$ $0.634\pm 0.016$ $0.338\pm 0.007$ $0.309\pm 0.006$ $0.264\pm 0.005$ $0.251\pm 0.005$ $0.265\pm 0.006$ $0.240\pm 0.005$ $0.211\pm 0.004$
MAE $0.602\pm 0.015$ $0.548\pm 0.012$ $0.382\pm 0.006$ $0.360\pm 0.006$ $0.320\pm 0.004$ $0.294\pm 0.004$ $0.285\pm 0.005$ $0.271\pm 0.004$ $0.239\pm 0.004$
ILI
MSE $4.480\pm 0.120$ $5.764\pm 0.150$ $3.483\pm 0.100$ $2.203\pm 0.070$ $1.572\pm 0.050$ $1.365\pm 0.045$ $0.952\pm 0.018$ $0.877\pm 0.016$ $0.672\pm 0.014$
MAE $1.444\pm 0.030$ $1.677\pm 0.035$ $1.287\pm 0.028$ $0.963\pm 0.022$ $0.891\pm 0.020$ $0.806\pm 0.018$ $0.793\pm 0.010$ $0.763\pm 0.010$ $0.505\pm 0.009$
Exchange Rate
MSE $0.968\pm 0.020$ $0.847\pm 0.018$ $0.197\pm 0.006$ $0.183\pm 0.005$ $0.175\pm 0.005$ $0.158\pm 0.004$ $0.146\pm 0.004$ $117\pm 0.003 0.0$$088\pm 0.003$
MAE $0.812\pm 0.015$ $0.752\pm 0.012$ $0.323\pm 0.006$ $0.297\pm 0.005$ $0.293\pm 0.004$ $0.281\pm 0.004$ $0.276\pm 0.005$ $0.258\pm 0.004$ $0.196\pm 0.004$
ETT
MSE $1.534\pm 0.030$ $1.410\pm 0.028$ $0.327\pm 0.007$ $0.305\pm 0.006$ $0.757\pm 0.015$ $0.291\pm 0.006$ $0.290\pm 0.006$ $0.275\pm 0.005$ $0.258\pm 0.005$
MAE $0.899\pm 0.020$ $0.810\pm 0.018$ $0.371\pm 0.006$ $0.349\pm 0.006$ $0.610\pm 0.012$ $0.333\pm 0.006$ $0.334\pm 0.006$ $0.323\pm 0.006$ $0.291\pm 0.005$